Given the following rates, what is the 2 year rate 1 year forward?

Years Rate(Annual
1 5%
2 5%
3 3%
Answer #1

Forward Rate of nth Year = (1+YTM of nth Year) to the power of n / (1+YTM of n-1 year) to the power of n-1) -1

The above data and calculation is depicted in an excel as below

The formula is as below

Forward Rate of nth Year = (1+YTM of nth Year)^n -1
(1+YTM of n-1 year)^n-1

Forward rate for the 2nd year rate for 1 year forward means whats the 1 year forward rate for Year 2.

Forward rate of 2nd year = (1+0.05)^2/ (1+0.05)^1 -1

Forward rate of 2nd year = 5.00%.